# Efficiently Training 7B LLM with 1 Million Sequence Length on 8 GPUs

### v1

@@ 2. Preliminary
Figure 3. The architecture of a typical LLM.
gpt_arch
@@ 2.1.1. Architecture
As shown in Figure 3, the architecture of a typical LLM comprises an input embedding layer, multiple decoder-only transformer layers, and a final classifier layer. The embedding layer converts input tokens into continuous representations. Each decoder-only transformer layer constitutes a multi-head self-attention module with causal mask, and an Feed-Forward Network (FFN) module containing Fully-Connected (FC) networks. The classifier layer takes the hidden states produced by the transformer layers as input, and generates a probability distribution over the vocabulary.
@@ 2.1.3. The Challenge of Huge Memory Requirement in Long Context Training
In the original form of self-attention, which is the most critical module in LLMs, input tokens are initially projected into queries, keys, and values. The queries and keys are multiplied and then softmaxed to create an $s\times s$ matrix that represents attention weights, which are subsequently used to compute a weighted sum of the values. Thus, it exhibits a memory complexity of $O(s^{2})$.
As the sequence length increases, storing the entire $s\times s$ tensors for the backward pass becomes infeasible. FlashAttention (Dao et al., 2022; Dao, 2023) has emerged as a solution for efficient attention computation in long context training. It minimizes memory I/O and carries out the attention computation in a streaming fashion in both the forward and backward pass, and thereby successfully get rid of the huge $s\times s$ matrix. With such designs, FlashAttention accelerates the overall computation speed (while maintaining a time complexity of $O(s^{2})$), and eliminates the $O(s^{2})$ memory requirement. Currently, FlashAttention has become the de-facto strategy for self-attention computation. Thus, we assume FlashAttention is employed throughout this work.
@@
2.2. Memory Reduction Techniques
Limited GPU memory has become the bottleneck of long context LLM training. To alleviate the memory pressure, there are two notable memory reduction techniques, namely activation recomputation and swapping.
Activation recomputation (Chen et al., 2016; Kirisame et al., 2021; Korthikanti et al., 2022) (a.k.a. activation checkpointing) selectively stores only the inputs of certain layers rather than storing all intermediate activations. During the backward pass, the required activations are recomputed on-the-fly. While this approach reduces the activation memory footprint required for LLM training, it introduces additional computation, which impacts efficiency. Swapping (Rhu et al., 2016; NVIDIA, 2024b; Ren et al., 2021), also known as CPU offloading, aims to relieve the GPU memory pressure by offloading GPU tensors to CPU memory, and fetch them back to GPU when needed. Although swapping does not consumes GPU computation units, it would still slow down the training if the data transmission cannot be overlapped by GPU computation. In general, both the two memory reduction techniques release the memory of activations in the forward pass, but need to rematerialize them in the backward pass, at the price of extra computation or data transmission overhead, respectively.
@@ 2.3. Distributed Parallelism Strategies
Distributed training is essential for efficiently training LLMs, especially in the senario of long context training. To facilitate the training of large-scale data and model, several distributed parallelism strategies have been proposed.
@@
Zero-Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2019) is a series of variants built upon DP, aiming to alleviate memory pressure. In naive DP, model parameters, gradients and optimizer states are replicated among all devices. ZeRO is designed in three stages to reduce these memory requirements respectively. First, ZeRO-1 partitions the optimizer states among all DP workers. Next, ZeRO-2 extends ZeRO-1 by also partitioning gradients, further reducing memory footprint. Finally, ZeRO-3, based on ZeRO-2, partitions model parameters among DP workers, further mitigating memory pressure but introducing additional communication to gather parameters during forward and backward pass.
@@
Pipeline Parallelism (PP) (Huang et al., 2019; Narayanan et al., 2021) is also proposed to address the problem that LLMs cannot be fit onto a single device. Different from TP, PP partitions different model layers into several stages, then distributes the stages to different devices. The input data is processed through these stages in a pipeline fashion. Given the peer-to-peer communication style, the PP stages are often distributed across nodes. However, PP introduces a phenomenon known as “bubble”, which corresponds to GPU idle time. The issue becomes more severe when the number of micro-batches is small.
To facilitate efficient long context training, several novel parallelism strategies have been recently proposed.
Sequence Parallelism (SP) (Korthikanti et al., 2022) is built upon TP to further reduce activation memory overhead. It splits the sequence dimension in the part of the model that does not apply TP. The original AllReduce communication now transitions to AllGather and ReduceScatter communication.
@@
Context Parallelism (CP) (Liu et al., 2023b; Li et al., 2023; NVIDIA, 2024a) proposes sharding the query, key, value matrices within the self-attention module along the sequence dimension across different devices. During attention computation, necessary communications are involved to ensure consistent results. The communications can be overlapped with computation by careful scheduling.
@@
Figure 4. An example memory request sequence of a transformer layer’s forward and backward propagation. Tensor 15 and 16 are skeletal tensors, while the others are transient tensors.
An example memory sequence
@@ 3. Anatomy and System Desiderata
Given the fact that the memory consumption of activations scales proportionally w.r.t. the sequence length in LLM training, we first provide an in-depth anatomy of the key characteristics of different activations in this section. Based on this analysis, we present the design desiderata that motivate the development of Memo.
@@ 3.1. Categorization of Activation Tensors
To be specific, according to their life cycles, we categorize activations generated during the forward propagation into two classes, which are the skeletal activations and the transient activations, where the former is necessary for the backward propagation while the latter is not.
For illustration, in Figure 4, tensors 13, 14, 17, 18, and 19 are produced during the forward propagation of a transformer layer, and are discarded before the completion of this layer’s forward pass. Similarly, tensors 20, 21, 22, 23, and 24 are generated during the backward propagation of this layer, and are discarded after corresponding computation. We term them “transient tensors” because they are both created and discarded within a single layer’s forward or backward pass. Transient tensors usually serve as temporary results. Conversely, tensors 15 and 16 are generated during the forward propagation and are needed for backward propagation, so they are discarded in this layer’s backward pass. We refer to these tensors as “skeletal tensors” because they are produced during the forward pass, and are essential for the gradient calculation during the backward pass.
@@ 3.2. Analysis of Skeletal Activations
Figure 6. Forward and backward propagation with rounding buffers for token-wise recomputation/swapping. During forward propagation, the darker part in the rounding buffers is offloaded to CPU, while the lighter part is discarded; during backward propagation, the darker part in the rounding buffers if prefetched from CPU, while the lighter part it recomputed.
@@
The most important characteristic of skeletal activations is that they are needed by backward computation, so they must reside in GPU memory at least before the backward propagation of the corresponding transformer layer begins. However, there is no doubt that maintaining all skeletal activations for backward propagation is infeasible. To this end, memory-saving techniques like activation recomputation and swapping become necessary for long context training. These techniques first release the skeletal activations of a transformer layer in the forward propagation, and later rematerialize them before the corresponding backward propagation. In essence, both the two techniques trade time for memory — the activation recomputation technique incurs extra computation overhead while the swapping technique necessitates transmitting the activations from CPU memory to GPU memory. For extremely long context lengths, using either technique would take significant time to rematerialize the skeletal activations, causing performance degradation.
@@
As a result, we desiderate a meticulous orchestration of the two memory-saving techniques to manage the skeletal activations, so that we can minimize the extra overhead while accommodating the huge memory requirement in long context training of LLMs. To achieve this goal, we develop a token-wise activation recomputation and swapping mechanism, which will be demonstrated in Section 4.1.
3.3. Analysis of Transient Activations
Transient activations are intermediate results generated and discarded during the forward (or backward) pass of a transformer layer. Actually, there are more transient activations than skeletal activations in a transformer layer. Specifically, we observe that the number of transient activations can exceed 5 times that of skeletal activations. Without careful management, the frequent allocation and deallocation can lead to memory fragmentation and degrades system performance.
Fortunately, the most important characteristic of transient activations is that they are identical cross transformer layers, which provides us with the opportunity to manage and reuse their memory regions to minimize fragmentation. In particular, the memory addresses of a single transformer layer’s transient activation tensors can be reused by all other transformer layer’s corresponding transient activation tensors. However, in practice, memory reuse is not fulfilled because the PyTorch caching allocator lack prior information of the memory request sequence during training iterations. This inspires us to statically plan the memory addresses of each transformer layer’s transient tensors, which will be described in detail in Section 4.2.
@@ 4. Memo Design
In this section, we propose Memo for fine-grained activation memory management. Our proposed method leverages fine-grained and structured activation management, akin to concise memos that share vital information. The main challenge of long context training is the large activation size which scales linearly w.r.t. sequence length. And we propose token-wise activation recomputation and swapping, along with a bi-level memory planning to address the issue, which targets skeletal activation tensors and transient activation tensors, respectively. The overview of Memo is depicted in Figure 2.
@@ 4.1. Token-wise Recomputation and Swapping
Skeletal tensors, generated during the forward pass of a transformer layer, must reside in GPU memory for the subsequent backward propagation. In practice, as sequence length grows, the size of skeletal activations increases linearly, which can easily exceed the capacity of GPU memory. As introduced in Section 2.2, currently the most widely-used technique to tackle this issue is activation recomputation, which stores only the input of each transformer layer, and discard the rest skeletal activation tensors of this layer. Prior to backward propagation of each layer, an additional forward pass of the layer is conducted to reconstruct all skeletal tensors so that the backward computation can be carried out. However, we note that the vanilla activation recomputation strategy is not an optimal choice to handle the challenge of linearly increasing skeletal activation memory, considering the following two reasons: (1) activation recomputation introduces redundant computation, thus diminishing training efficiency; and (2) the memory overhead of retaining the input tensor of each transformer layer can still be expensive, especially when the sequence length is too long or the number of layers is too large. Take the training of GPT-7B with a context length of 1 million as an example again. For only one sequence, the input tensors of all 32 transformer layers together consume 128GB. Even using a SP degree of 8, it takes 16GB for each GPU to store the input tensors of all 32 transformer layers, which already takes up to 20% of total GPU memory capacity.
@@
For backward propagation, after the backward pass of transformer layer $(i+2)$ ends, the contents within rounding buffer $(i\%2)$ become useless, and we start prefetching the skeletal activations of transformer layer $i$ to rounding buffer $(i\%2)$ using another CUDA stream. The prefetching of transformer layer $i$’s skeletal activations happens simultaneously with the backward propagation of transformer layer $(i+1)$. When the sequence length is sufficiently long, with careful computation-transmission overlapping and synchronization, CPU swapping can substitute activation recomputation without incurring additional computation overhead.
@@
  • For sequence lengths that are not sufficiently long, the time required to offload all skeletal activations to CPU memory surpasses the computation time for a single transformer layer. This discrepancy forces the computation of transformer layer $(i+2)$ to be delayed until the offloading of rounding buffer to CPU memory is completed, thereby blocks the normal GPU computation workflow. For instance, as illustrated in Figure 1(b), when training a 7B GPT model on 8 GPUs with a TP size of 8, ideal overlap between the computation of a transformer layer and the offloading of its skeletal activations occurs only for sequence lengths exceeding 192K. In practice, the sequence lengths of most training datasets for LLMs are moderate and may be not sufficient to ensure an ideal overlap between computation and transmission.
  • In theory, a longer sequence length provides more opportunities for overlapping CPU offloading with GPU computation. However, in practice, the CPU memory in GPU servers is often limited. For a typical GPU server which has several terabytes CPU memory (e.g. 2TB in our environment), this is insufficient to store all skeletal activations when the sequence length is excessively long or the number of transformer layers is too large. For instance, when training the 7B model on a server equipped with 8 GPUs using a sequence length of 1 million, the skeletal activations amount to a total size of 4096GB, which is double the capacity of CPU memory.
Figure 7. Forward time of FlashAttention and other parts of a transformer layer when training a 7B GPT on 8 GPUs with a TP size of 8.
@@
Therefore, instead of simply offloading all skeletal activations to CPU memory, we employ selective activation swapping to ensure perfect overlap of computation and transmission for short sequences as well as to avoid depleting CPU memory for extremely long context lengths. Memo manages to determine the selection of swapping at both the tensor and token granularities, as depicted in Figure 6.
@@
At the token granularity, we develop the token-wise activation recomputation and swapping technique to reduce the memory consumption of all skeletal activation tensors other than the output of FlashAttention and the input of each transformer layer. To be specific, as shown in Figure 6, for each of these skeletal activation tensors, we only offload a fraction (denoted as $\alpha$) to CPU, while the remaining part is discarded, ensuring perfect overlapping and to avoid CPU OOM error. Before the backward propagation, the discarded part is rematerialized via recomputation while the offloaded part is prefetched.
@@
(1) $\max\quad$ $\alpha,$
(2) s.t. $(S_{input}+S_{attn}+\alpha\cdot S_{others})/B\leq T_{layer},$
(3) $(n-2)(S_{input}+S_{attn}+\alpha\cdot S_{others})\leq M_{CPU}.$
where $S_{input}$, $S_{attn}$, and $S_{other}$ stand for the size of input tensor, the size of FlashAttention output tensor, the total size of other skeletal activation tensors, respectively, $B$ is the PCIe bandwidth between GPU and CPU, $T_{layer}$ is the forward time of a single transformer layer, $n$ is the total number of transformer layers, and $M_{CPU}$ stands for the capacity of CPU memory. It is worth noting that, the last two transformer layers can initiate the backward pass immediately after the forward pass, obviating the need for swapping. These variables can be easily obtained through profiling before training, so we can determine an appropriate $\alpha$ without much effort.
As a special case, when the determined fraction is 0, skeletal buffers except layer input and attention output can share one GPU buffer, as they are fully recomputed and do not need to be offloaded. Thus there is no need to use two rounding buffers to avoid data corruption.

### v2

@@ 2. Preliminary
(a) The architecture of a typical LLM.
(b) An example memory request sequence.
Figure 3. (a): The architecture of a typical LLM. (b): An example memory request sequence of a transformer layer’s forward and backward pass. Tensors 15 and 16 are skeletal tensors, while the others are transient tensors.
@@ 2.1.1. Architecture
As shown in Figure 3(a), the architecture of an LLM comprises an input embedding layer, multiple transformer layers, and a final classifier layer. The embedding layer converts input tokens into continuous representations. Each decoder-only transformer layer constitutes a multi-head self-attention module with causal mask, and an Feed-Forward Network (FFN) module containing Fully-Connected (FC) networks. The classifier layer takes the hidden states produced by the transformer layers as input, and generates a probability distribution over the vocabulary.
@@ 2.1.3. The Challenge of Huge Memory Requirement in Long Context Training
Self-attention is the most critical module in LLMs. It facilitates information interaction between tokens: the input tensor is first projected into query ($Q$), key ($K$), and value ($V$), each with the shape ($b,s,h$); then the output is given by $O=\text{softmax}({QK^{T}}/{\sqrt{d}})\cdot V$, which incurs $O(s^{2})$ time and space complexity due to the ($b,s,s$) matrices. FlashAttention (Dao et al., 2022; Dao, 2024), the de-facto attention implementation in nowadays LLM computation, processes the computation in tiles, discards intermediate results, and maintains compact states to generate the final output $O$. This method avoids storing the $O(s^{2})$ matrices. During the backward pass, FlashAttention re-computes these intermediate results in a tiled manner for gradient calculation. Thanks to this design, FlashAttention significantly reduces memory requirements to just $O(s)$ complexity. Although several alternative approaches like sparsification (Miao et al., 2024b) (e.g., BigBird (Zaheer et al., 2020), Longformer (Beltagy et al., 2020)), kernelization (e.g., LinearAttn (Katharopoulos et al., 2020), CosFormer (Qin et al., 2022)), and low-rank approximation (e.g., Linformer (Wang et al., 2020), Nyströmformer (Xiong et al., 2021)) also aim to reduce the quadratic memory demands of self-attention, these methods modify the attention mechanism and can potentially compromise accuracy. In this paper, we adopt FlashAttention as the default method, considering its prevalent use in practice.
@@
$2.$ Memory Reduction Techniques
As mentioned in Section 1, when existing query processing engines face the challenge of limited high-bandwidth memory, a common strategy to mitigate memory pressure involves discarding certain data structures and rematerializing them as needed (Zaharia et al., 2016; Johnson et al., 2021; Phan and Li, 2008). There are two prevalent techniques: (1) recomputing the results and (2) swapping data to lower-tier memory and retrieving it when necessary. Both methods are also widely-used in neural network training to rematerialize activation tensors.
Activation recomputation (Chen et al., 2016; Kirisame et al., 2021; Korthikanti et al., 2022) (a.k.a. activation checkpointing) selectively stores the inputs of certain layers rather than all intermediate activations. During the backward pass, the required activations are recomputed on-the-fly. While this approach reduces the activation memory footprint required for LLM training, it introduces additional computation, which impacts efficiency. Swapping (Rhu et al., 2016; NVIDIA, 2024b; Ren et al., 2021), also known as CPU offloading, aims to relieve the GPU memory pressure by offloading GPU tensors to CPU memory, and fetch them back to GPU when needed. Through careful scheduling, the data transmission overhead can be overlapped with GPU computation, a technique also popular in GPU databases (Johnson et al., 2021; Rui et al., 2020; Wang et al., 2014). However, if data transmission is too time-consuming to overlap, swapping can significantly slow down training. In general, both memory reduction techniques release the memory of activations in the forward pass, but need to rematerialize them in the backward pass, at the price of extra computation or data transmission overhead, respectively.
@@ 2.3. Distributed Parallelism Strategies
Distributed training is essential for efficiently training LLMs, especially in scenarios of long context training. To facilitate the training of large-scale data and model, several distributed parallelism strategies have been proposed.
@@
Zero-Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2019) is a series of variants built upon DP, aiming to alleviate memory pressure. Naive DP replicates model parameters, gradients and optimizer states among all devices. ZeRO is designed in three stages to reduce these memory requirements respectively. First, ZeRO-1 partitions the optimizer states among all DP workers. Next, ZeRO-2 extends ZeRO-1 by also partitioning gradients, further reducing memory footprint. Finally, ZeRO-3, based on ZeRO-2, partitions model parameters among DP workers, further mitigating memory pressure but introducing additional communication to gather parameters during training.
@@
Pipeline Parallelism ($P$) (Huang et al., 2019; Narayanan et al., 2021; Guan et al., 2024) is also proposed to address the problem that LLMs cannot be fit into a single device. Different from TP, $P$ partitions model layers into several stages, then distributes the stages to different devices. The input data is processed through these stages in a pipeline fashion. Given the peer-to-peer communication style, the $P$ stages are often distributed across nodes. However, $P$ introduces a phenomenon known as “bubble”, which corresponds to GPU idle time. The issue becomes more severe when the number of micro-batches is small.
To facilitate efficient long context training, several novel parallelism strategies have been proposed recently.
Sequence Parallelism (SP) (Korthikanti et al., 2022) is built upon TP to further reduce activation memory overhead. It splits the sequence dimension in the part of the model that does not apply TP. The original AllReduce communication now transitions to AllGather and ReduceScatter.
@@
Context Parallelism (CP) (Liu et al., 2023; Li et al., 2023a; NVIDIA, 2024a) shards the query, key, and value matrices within the attention module along the sequence dimension across different devices. During attention computation, necessary communications are involved to ensure consistent results, which can be overlapped with computation by careful scheduling.
@@ 3. Anatomy and System Desiderata
Managing fragmented massive data storage is a critical data management issue, particularly when workloads are constrained by limited high-bandwidth memory. Long-context LLM training demonstrates such a challenge, where huge fragmented activation memory significantly impedes efficient training within constrained GPU memory resources.
In this section, we first provide an in-depth anatomy of the key characteristics of activation data storage in long-context LLM training. Based on this analysis, we present the design desiderata that motivates the development of Memo.
@@ 3.1. Categorization of Activation Tensors
In long-context LLM training, the primary memory consumption originates from activation tensors, which are the outputs of computing operators in LLMs. According to their life cycles, we can categorize activations generated during the forward propagation into two classes, which are the skeletal activations and the transient activations, where the former is necessary for the backward propagation while the latter is not.
For illustration, in Figure 3(b), tensors 13, 14, 17, 18, and 19 are produced during the forward pass of a transformer layer, and are discarded before the completion of this layer’s forward pass. Similarly, tensors 20, 21, 22, 23, and 24 are generated during the backward pass of this layer, and are discarded after corresponding computation. We term them “transient tensors” because they are created and discarded within a single layer’s forward or backward pass. Transient tensors usually serve as temporary results. Conversely, tensors 15 and 16 are generated during the forward propagation and are needed for backward propagation, so they are discarded in this layer’s backward pass. We refer to these tensors as “skeletal tensors” because they are produced during the forward pass, and are essential for the gradient calculation during the backward pass.
@@ 3.2. Analysis of Skeletal Activations
Figure 5. Forward and backward propagation with rounding buffers for token-wise recomputation/swapping. During forward propagation, the darker part in the rounding buffers is offloaded to CPU, while the lighter part is discarded; during backward propagation, the darker part in the rounding buffers is prefetched from CPU, while the lighter part is recomputed.
@@
An important characteristic of skeletal activations is that they are needed by backward computation, so they must reside in GPU at least before the backward propagation of the corresponding transformer layer begins. However, maintaining all skeletal activations for backward propagation is infeasible. To this end, memory-saving techniques like recomputation and swapping, which are also prevalent in traditional data management problems (Zaharia et al., 2016; Johnson et al., 2021; Phan and Li, 2008), become necessary for long context training.
In LLM training, these techniques first release the skeletal activations of a transformer layer in the forward propagation, and later rematerialize them before the corresponding backward propagation.
@@
Unfortunately, naïvely applying activation recomputation or swapping is insufficient to tackle the challenge of managing large-scale skeletal activations. Both techniques trade time for memory — the activation recomputation technique incurs extra computation overhead while the swapping technique necessitates transmitting the activations from CPU memory to GPU memory. Using recomputation alone incurs significant additional computation overhead, while employing swapping alone can lead to CPU OOM error (when the sequence is too long) or block GPU computation (when the swapping time cannot be fully overlapped). As a result, we desiderate a meticulous orchestration of the two memory-saving techniques to manage the skeletal activations, so that we can minimize the extra overhead while accommodating the huge memory requirement in long context training of LLMs. To achieve this, we develop a token-wise activation recomputation and swapping mechanism, which will be demonstrated in Section 4.1.
$3.$ Analysis of Transient Activations
Transient activations are intermediate results generated and discarded during the forward (or backward) pass of a transformer layer. Actually, there are more transient activations than skeletal activations in a transformer layer. Specifically, we observe that the number of transient activations can exceed 5 times that of skeletal activations. Without careful management, the frequent allocation and deallocation can lead to memory fragmentation, which degrades system performance and poses a major concern for managing massive data storage. There have been studies in the data management community focusing on memory defragmentation (Durner et al., 2019; Oukid et al., 2017), leveraging the characteristics of target workloads to devise appropriate and innovative methods. Following the methodology, we analyze the LLM training process and attempt to defragment the tensor memory. During training, memory requests are identical across both transformer layers and training iterations, providing an opportunity to manage and reuse these memory regions effectively to minimize fragmentation. In particular, the memory addresses of a single transformer layer’s transient activation tensors can be reused by all other transformer layer’s corresponding transient activation tensors. However, in practice, memory reuse is not fulfilled because the PyTorch caching allocator lack prior information of the memory request sequence during training iterations. This inspires us to statically plan the memory addresses of each transformer layer’s transient tensors, which will be described in detail in Section 4.2.
@@ 4. Memo Design
In this section, we propose Memo for fine-grained activation memory management. Our proposed method leverages fine-grained and structured activation management, akin to concise memos that share vital information. The main challenge of long context training is the large activation size which scales linearly w.r.t. sequence length. We propose token-wise activation recomputation and swapping, along with a bi-level memory planning to address the issue, which targets skeletal activations and transient activations, respectively. The overview of Memo is depicted in Figure 2.
@@ 4.1. Token-wise Recomputation and Swapping
Skeletal tensors, generated during the forward pass of a transformer layer, must reside in GPU memory for the subsequent backward propagation. In practice, as sequence length grows, the size of skeletal activations increases linearly, which can easily exceed the capacity of GPU memory. As introduced in Section 2.2, currently the most widely-used technique to tackle this issue is activation recomputation, which stores only the input of each transformer layer, and discards the rest skeletal activation tensors of this layer. Prior to backward propagation of each layer, an additional forward pass of the layer is conducted to reconstruct all skeletal tensors so that the backward computation can be carried out. However, we note that the vanilla activation recomputation strategy is not an optimal choice to handle the challenge of linearly increasing skeletal activation memory, considering the following two reasons: (1) activation recomputation introduces redundant computation, thus diminishing training efficiency; and (2) the memory overhead of retaining the input tensor of each transformer layer can still be expensive, especially when the sequence length is too long or the number of layers is too large. Take the training of GPT-7B with a context length of 1 million as an example again. For only one sequence, the input tensors of all 32 transformer layers together consume 128GB. Even using a SP degree of 8, it takes 16GB for each GPU to store the input tensors of all 32 transformer layers, which already takes up to 20% of total GPU memory capacity.
@@
For backward propagation, after the backward pass of transformer layer $(i+2)$ ends, the contents within rounding buffer $(i\%2)$ become useless, and we start prefetching the skeletal activations of transformer layer $i$ to rounding buffer $(i\%2)$ using another CUDA stream. The prefetching of transformer layer $i$’s skeletal activations happens simultaneously with the backward propagation of transformer layer $(i+1)$. When the sequence length is sufficiently long, with careful computation-transmission overlapping and synchronization, CPU swapping can substitute activation recomputation without incurring additional overhead.
@@
  • For sequence lengths that are not sufficiently long, the time required to offload all skeletal activations to CPU memory surpasses the computation time for a single transformer layer. This discrepancy forces the computation of transformer layer $(i+2)$ to be delayed until the offloading of rounding buffer to CPU memory is completed, thereby blocks the normal GPU computation workflow. For instance, as illustrated in Figure 1, when training a 7B GPT model on 8 GPUs with a TP degree of 8, ideal overlap between the computation of a transformer layer and the offloading of its skeletal activations occurs only for sequence lengths exceeding 192K. In practice, the sequence lengths of most training datasets are moderate and may be not sufficient to ensure an ideal overlap between computation and transmission.
  • In theory, a longer sequence length provides more opportunities for overlapping swapping with GPU computation. However, in practice, the CPU memory is often limited. For a typical GPU server which has several terabytes CPU memory (e.g. 2TB in our environment), it is insufficient to store all skeletal activations when the sequence length is excessively long or the number of transformer layers is too large. For instance, when training the 7B model on a server equipped with 8 GPUs using a sequence length of 1 million, the skeletal activations amount to a total size of 4096GB, which is double the capacity of CPU memory.
Figure 6. Forward time of FlashAttention and other parts of a transformer layer when training a 7B GPT on 8 GPUs with a TP degree of 8.
@@
Therefore, instead of simply offloading all skeletal activations to CPU memory, we employ selective activation swapping to ensure perfect overlap of computation and transmission for short sequences as well as to avoid depleting CPU memory for extremely long context lengths. Memo determines the selection of swapping at both the tensor and token granularities, as depicted in Figure 5.
@@
At the token granularity, we develop the token-wise activation recomputation and swapping technique to reduce the memory consumption of all skeletal activation tensors other than the output of FlashAttention and the input of each layer. To be specific, as shown in Figure 5, for each of these skeletal activation tensors, we offload a fraction (denoted as $\alpha$) to CPU, while the remaining part is discarded, ensuring perfect overlapping and to avoid CPU OOM error. Before the backward pass, the discarded part is rematerialized via recomputation while the offloaded part is prefetched.
@@
$\max\quad$ $\alpha,$
s.t. $(S_{input}+S_{attn}+\alpha\cdot S_{others})/B\leq T_{layer},$
$(n-2)(S_{input}+S_{attn}+\alpha\cdot S_{others})\leq M_{CPU}.$
where $S_{input}$, $S_{attn}$, and $S_{other}$ stand for the size of input tensor, the size of FlashAttention output tensor, the total size of other skeletal activation tensors, respectively; $B$ is the PCIe bandwidth between GPU and CPU, $T_{layer}$ is the forward time of a single transformer layer, $n$ is the total number of transformer layers, and $M_{CPU}$ stands for the capacity of CPU memory. It is worth noting that, the last two transformer layers can initiate the backward pass immediately after the forward pass, obviating the need for swapping. These variables can be easily obtained through profiling before training, so we can determine an appropriate $\alpha$ without much effort.